OpenStudy (anonymous):

Explain quantitatively the relationship between linear and rotational motion.

OpenStudy (matthewrlee):

How about a please?

OpenStudy (anonymous):

A force of F=MA excert on a object. Then the object will have a average velocity of : Delta position/Delta Time While rotational velocity is the same but will be: 2(pi)r*Angle/360/Delta Time

OpenStudy (anonymous):

Thanks for the reply, I copied the question directly from an assignment sorry if I sounded abrupt.

OpenStudy (anonymous):

\[s = r \theta\] \[v = r \omega\] \[a = r \alpha\] where r is the radius, s is the arc-length, a is the tangential acceleration, v is the tangential velocity, theta as the angle measured in radians, omega is the angular velocity measured in radians/s and alpha is the angular acceleration measured in radians per second per second

OpenStudy (anonymous):

Ok, thanks for the answers guys. Am I right in thinking that Stan gave a much more fitting answer to the question or am I missing something?

OpenStudy (anonymous):

the equations of motions are true for circular motion. replace v with \[\omega\] u with \[\omega\](i) a with \[\alpha\] s with \[\Theta\]